OpenStudy (anonymous):

Calculate the average rate of change for the graphed sequence from n = 1 to n = 5.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (calculusxy):

Isn't finding the rate of change also known as the slope? Or like putting it into slope-intercept form?

OpenStudy (michele_laino):

the requested rate, is given by the subsequent formula: \[\Large r = \frac{{{a_5} - {a_1}}}{{5 - 1}} = \frac{{14 - 2}}{{5 - 1}} = ...?\]

OpenStudy (anonymous):

12/4

OpenStudy (anonymous):

3

OpenStudy (calculusxy):

yes that's your slope

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Thanks:)!!!!

OpenStudy (michele_laino):

:) :)